OpenStudy (anonymous):

What is the integer solution of the following equation e^(2x)-3e^(x)=-2

OpenStudy (dumbcow):

solve like a quadratic u^2 -3u +2 = 0 where u = e^x

OpenStudy (ranga):

Yes. Use the substitution suggested above. But if you wish to you can directly factor it by grouping: e^(2x) - 3e^x = -2 e^(2x) - 3e^x + 2 = 0 e^(2x) - e^x - 2e^x + 2 = 0 e^x(e^x - 1) - 2(e^x - 1) = 0 (e^x - 1)(e^x - 2) = 0 e^x = 1 or e^x = 2 solve for x.

OpenStudy (anonymous):

Is there a reason that when you get to the 4th step there is no 3?

OpenStudy (ranga):

On line 3, the -3e^x is split into -e^x - 2e^x to help with the factoring. Was that your question?

OpenStudy (anonymous):

yes that was my question. thanks:) when solving for x, I put log on both sides?

OpenStudy (ranga):

yes. you take natural log or ln on both sides. Remember log or ln of 1 is 0 and ln(e) = 1

OpenStudy (anonymous):

really lost now…

OpenStudy (ranga):

e^x = 1 or e^x = 2 start with e^x = 1 take ln on both sides: ln(e^x) = ln(1) x * ln(e) = ln(1) x * 1 = 0 x = 0 e^x = 2 ln(e^x) = ln(2) x * ln(e) = ln(2) x * 1 = ln(2) x = ln(2) x = 0 or ln(2)

OpenStudy (ranga):

\[\Large \ln(a)^m = m * \ln(a)\]

OpenStudy (ranga):

\[\Large \ln(e)^x = x * \ln(e) = x\]

OpenStudy (anonymous):

the first one is the equation and the second one has everything plugged in?ln(a)m=m∗ln(a) ln(e)x=x∗ln(e)=x

OpenStudy (ranga):

The first one is an IDENTITY. That means it is good for a whole range of a and m. Whenever we come across ln(something)^exponent the identity says it can be written as: ln(something)^exponent = exponent * ln(something) In the second case we make use of the identity to rewrite ln(e^x) as x * ln(e) and ln(e) = 1.

OpenStudy (ranga):

Examples:\[\Large \ln(5)^8 = 8 * \ln(5)\]\[\Large \ln(e)^5 = 5 * \ln(e) = 5\]

OpenStudy (anonymous):

Thanks you very much :) am beginning to understand it more. which is a positive considering i have a midterm tomorrow.

OpenStudy (ranga):

cool. There are just a few logarithmic identities. And once you familiarize yourself you can breeze through problems involving logarithms.

OpenStudy (ranga):

you are welcome.